OpenStudy (anonymous):

Can someone Explain this step-By-step? Can someone Explain step-by-step how to solve this? Find the equation of the perpendicular line to each of the following that passes through the point indicated. 4x + 5y = - 40 passing through the point (-1, 3)

OpenStudy (anonymous):



OpenStudy (anonymous):

so, the slope you need is 5/4

OpenStudy (anonymous):

-5x + 4y = 17 5x + 4y = -17 5x - 4y = -17 5x + 4y = 17 these are the answers that i have the choice of getting

OpenStudy (anonymous):